Question

An inspector reported a total of 10 defects out of a total batch size of 200 parts. On closer examination, it was determined that 4 defective parts were mistakenly placed in the batch of acceptable items. What is the inspector’s accuracy in this instance? Specifically, what are the values of p1 and p2. (b) What was the true fraction defect rate q?
Answer #1

(a)

Actual number of defective units =10+4 = 14

Actual number of good units = 200 - 14 = 186

Number of good units correctly detected by the Inspector = 200-10-4 = 186

Number of defective units correctly detected by the Inspector = 10

-

p1 = Number of good units correctly detected by the Inspector / Actual number of good units

= 186/186

= 1

-

p2 = Number of defective units correctly detected by the Inspector / Actual number of defective units

= 10/14

= 0.7143

--------------------------------------------------------------------

(b)

True fraction defect rate, q = Actual number of defective units / Total batch size

= 14/200

= 0.07